OpenStudy (anonymous):

The 11 th term and 29 th term of an AP. are in the ratio 3: 8. What is the ratio of sum of first 13 terms and first 17 terms of the AP

OpenStudy (anonymous):

th term of an AP is a+(n-1)d where a is first term ,n are no. of terms and d is the common differnce

OpenStudy (anonymous):

a+10d/a+28d=3/8

OpenStudy (anonymous):

8a+80d =3a+84d 5a=4d a=4/5d

OpenStudy (anonymous):

sum of first 13 terms and first 17 terms of the AP.. can be written as {13/2[2a+12d]}/{17/2[2a+16d]} here , i have to put a value

OpenStudy (anonymous):

tell me i have done correct or not

OpenStudy (anonymous):

some budy plse tell me if i have done right or wrong